OpenStudy (zootedyoungin):

two coins are tossed what is the theoretical probability that one lands on heads and the other lands on tails

OpenStudy (zootedyoungin):

@agent0smith

OpenStudy (agent0smith):

What is the probability of a coin landing on heads? Same question for tails? Multiply.

OpenStudy (p0sitr0n):

P(A and B)=P(A)P(B) IF they are independent

OpenStudy (zootedyoungin):

wo that is confusing

OpenStudy (zootedyoungin):

would it be 50% 1/2?

OpenStudy (agent0smith):

What is the probability of a coin landing on heads? Same question for tails? Multiply.

OpenStudy (zootedyoungin):

or same as the others because there is two coins 1/4 0.25

OpenStudy (zootedyoungin):

Heads = 1/2 (50% Tails = 1/2 (50%) 0.5x0.5 = 0.25 Ans=1/4?